OpenStudy (anonymous):

point a is at 3r point b is at 2r from q determine elecric potential Vab in term of kq/r?

OpenStudy (anonymous):



OpenStudy (ash2326):

Vab=Va-Vb We know potential is given by this relation at a distance r from a charge Q \[V=\frac{KQ}{r}\] \[V_a=\frac{KQ}{3r}\] \[V_b=\frac{KQ}{2r}\] \[Va< Vb\] so \[V_{ab}<0\] Could you do it now?

OpenStudy (anonymous):

i have solved but its -kq/6.where answer is required in r form

OpenStudy (anonymous):

Vab=kq(1/2r-1/3r) Vab=kq(3r-2r)/6r =kqr/6r =kq/6 I have solved like this

OpenStudy (ash2326):

\[V_{ab}=kq(\frac{1}{3r}-\frac{1}{2r})\] LCM of 2r and 3r is 6r \[V_{ab}=kq\frac{2-3}{6r}\] \[V_{ab}=-\frac{kq}{6r}\]

OpenStudy (anonymous):

thanks you I was wrong in solving.

OpenStudy (ash2326):

:D Welcome